OpenStudy (anonymous):

Given that f(x)=x-1 and g(x)=x^2-4, evaluate gog at x=2

OpenStudy (anonymous):

\[g(2)=2^2-4=0\] \[g(g(2))=f(0)=0^2-4=-4\]

OpenStudy (anonymous):

last line should have read \[g(g(2))=g(0)=0^2-4=-4\]

OpenStudy (anonymous):

Thanks so much!